Find the force required to make a 50 kg mass of metal with a S.G of 5.5,when it is half submerged in a. water b. Oil with S.G of 0.68?
1
Expert's answer
2020-03-16T13:20:24-0400

a)


"W=F+0.5\rho gV"

"F_w=50(9.8)-0.5\frac{50(9.8)}{5.5}=445\ N"


b)


"F_o=50(9.8)-0.5\frac{50(0.68)(9.8)}{5.5}=460\ N"